A chemist wants to mix some 34% alloy with some 70% alloy to produce 30 ounces of a 40% alloy. How many ounces of each strength should he use?

Answer by stanbon(75887)

A chemist wants to mix some 34% alloy with some 70% alloy to produce 30 ounces of a 40% alloy. How many ounces of each strength should he use?
--------
Equation:
active + active = active
0.34x + 0.70(30-x) = 0.40*30
-----
34x + 70*30-70x = 40*30
-36x = -30*30
x = 25 ounces (amt. of 34% alloy needed)
------
70-25 = 55 ounces (amt. of 70% alloy needed)
================
Cheers,
Stan H.